# SmartGSN: a generative AI-powered online tool for the management of assurance cases

### v1

@@ III-A3 LLM settings
Our experiment focuses on two LLMs: GPT-4o and GPT-4 Turbo. To evaluate the generative AI-based pattern detection feature of SmartGSN, we performed that experiment five times (K=5) to account for the non-deterministic nature of the LLMs at hand. We set each LLM at hand to its default parameters. Thus, we set the Temperature to the default value of 1. Additionally, we set the maximum number of each LLM tokens to be generated at 4096.
III-A4 Metrics used to craft pattern detection prompts
@@
In the context of this paper, we focus on a set of two well-known metrics (i.e. n=2) that allow computing text similarity. These metrics are: BLEU Score and Semantic Similarity [16, 17]. To support pattern detection, SmartGSN uses these metrics to compute the similarity between the assurance case pattern and the assurance case, both formalized in an advanced structured prose format we introduced in previous work [14]. SmartGSN computes these two metrics by relying on the following Python libraries: Sacrebleu, and scikit-learn. For brevity’s sake, we will only report the experiment results we obtained with the following metric thresholds: 0.4, 0.6, and 0.8.
III-A5 Experiments description
@@
We rely on a combination of two prompting strategies to perform our experiment [18, 19]: Zero-shot + CoT. In these techniques, we do not include any examples in our prompt. Still, we provide a series of intermediate logical steps needed to improve our tool’s ability to detect a given pattern within an assurance case. Thus, to validate the pattern detection feature of SmartGSN, we used SmartGSN to perform an experiment in which we run each LLM five times with the Zero-shot prompting technique together with CoT to detect patterns in assurance cases.
III-A6 Metrics used to assess experiment results
To assess the results, we relied on well-known metrics: Precision, Recall, and F1-Measure.
III-B Results analysis: using SmartGSN for pattern detection
System Mod Threshold
0.4 0.6 0.8
R P F1 R P F1 R P F1
ACAS XU 4o 1 1 1 1 1 1 0 0 0
4T 1 1 1 1 1 1 0 0 0
BlueROV2 4o 0.5 1 0.67 0.5 1 0.67 0 0 0
4T 0.5 1 0.67 0.5 1 0.67 0 0 0
GPCA 4o 1 1 1 0 0 0 0 0 0
4T 1 1 1 0 0 0 0 0 0
IMSoftware 4o 1 1 1 1 1 1 0 0 0
4T 1 1 1 1 1 1 0 0 0
DeepMind 4o 1 1 1 0 0 0 0 0 0
4T 1 1 1 0 0 0 0 0 0
TABLE I: Recall (R), Precision (P) and F1-Score (F1) results
@@ III-B Results analysis: using SmartGSN for pattern detection
Table I reports the average of recall (R), precision (P), and F1-Measure (F1) for the five systems in our dataset using two models across various thresholds. We compute each average over five runs i.e. for K=5. The ”Mod” header either denotes GPT-4o (i.e. ”4o”) or GPT-4 Turbo (i.e. ”4T”). For each system, both models achieve perfect scores at threshold of 0.4, with the exception of BLUEROV2, which has an average recall of 0.5 and an average F1-Measure of 0.67. At a threshold of 0.6, both ACAS XU and IM Software maintain perfect scores, whereas the BLUEROV2 still yields the same results. The DeepMind and GPCA metric scores drops to zero at this threshold. At a threshold of 0.8, all systems and models achieve zero scores across all metrics, indicating an inability for SmartGSN to detect patterns within an assurance case at this threshold. Threshold values affect the pattern detection process; lower thresholds like 0.4 results in detecting more patterns, while at higher thresholds (e.g., 0.8), performance drastically declines. Hence, the optimal threshold range for effective detection of patterns within assurance cases in our dataset likely falls within the [0.4, 0.6[ interval. More experiments will help confirm that observation.
IV Conclusion and future work
@@
In this paper, we presented a novel tool called SmartGSN. The latter leverages large language models (LLMs) to support the semi-automatic management of assurance cases.
@@
Future work will focus on refining the pattern detection approach our tool supports by using more advanced heuristics. Future work will also focus on assurance case refactoring using assurance case patterns. In addition, future work will focus on implementing a collaboration feature that will support the concurrent edition of the same assurance case when using SmartGSN. We are also planning to carry out a user study to gauge the usability of our tool in various application domains.
Acknowledgment
We thank Kimya Khakzad Shahandashti for suggesting the use of React Flow for node-based UIs implementation.

### v2

@@ 4.1.3 Settings (customization) of the LLMs
Our experiment focuses on two well-known LLMs: GPT-4o and GPT-4 Turbo. To run experiments, we set each of these LLMs to its default parameters. Thus, for each LLM, we set the Temperature to the default value of 1. Additionally, we set the default maximum number of each LLM output tokens to be generated at 4096.
4.1.4 Metrics used to craft pattern detection prompts
In the context of this experiment, our metric-based rule focuses on a set of two well-known metrics (i.e. n=2) that allow computing text similarity. These metrics are:
  • BLEU Score: it is a widely used metric in Natural Language Processing and is one of the most common evaluation metrics for natural language texts. SmartGSN uses that metric to assess the lexical similarity (word-by-word similarity) between an assurance case pattern and an assurance case. The BLEU score varies between 0 and 1, where 0 indicates no match between the pattern(s) and the assurance case, and 1 represents a perfect match.
  • Cosine Similarity: SmartGSN uses this metric to assess the semantic similarity between texts in the GSN elements of an assurance case pattern and those in an assurance case. The cosine similarity score ranges from -1 to 1, where -1 indicates completely dissimilar texts and 1 signifies identical texts.
@@ 4.1.4 Metrics used to craft pattern detection prompts
To support pattern detection, SmartGSN uses these two metrics to compute the similarity between the assurance case pattern and the assurance case, both formalized in an advanced structured prose format we introduced in previous work [19]. SmartGSN computes these two metrics by relying on the following Python libraries: Sacrebleu, and scikit-learn. Each SmartGSN LLM then uses the values of both metrics to apply our metric-based rule when detecting patterns. Each metric threshold ranges from 0 to 1. For brevity’s sake, we will only report the experiment results we obtained with the following metric thresholds: 0.2, 0.4, 0.6, 0.8, and 1. In our experiment, we consider $threshold\_metric\_1$ and $threshold\_metric\_2$ are identical and are respectively associated with the BLEU score and the cosine similarity.
4.1.5 Prompt description
@@
We rely on a combination of two well-known prompting strategies to perform our experiment [28], [4]: Zero-shot + CoT (Chain-of-Thought). When combining both techniques, we do not therefore include any examples in each LLM prompt. Still, we provide a series of intermediate logical steps needed to improve our tool’s ability to detect a given pattern within an assurance case. Figures 4 and 5 illustrate the structure of our system prompt and user prompt, respectively, and can be accessed online  [^1] https://github.com/Oluwafemi17/SmartGSN/blob/109ab8ed180580435442a31ee458b25aafe3cced/System_Prompt.png  [^2] https://github.com/Oluwafemi17/SmartGSN/blob/109ab8ed180580435442a31ee458b25aafe3cced/User_Prompt.png.
Our LLM system prompt consists of the following elements: a sequence of intermediate steps to identify whether an assurance case pattern is present within an assurance case; context information outlining the various elements and decorators in an assurance case and an assurance case pattern represented in GSN; predicate-based rules for translating both patterns and assurance cases into a formalized format, enhancing the LLMs’ ability to understand the relationships among GSN elements in our dataset; and finally, domain information detailing the specialized knowledge and facts relevant to the system for which an assurance case pattern is to be detected. In previous work [19], we explained how we specified context information, derived predicate-based rules and collected domain information. Our LLM user prompt includes both the formalized assurance case pattern and the formalized assurance case from which the pattern needs to be detected. Consequently, each SmartGSN LLM relies on both system and user prompts to accurately detect a given pattern within a specific assurance case.
4.1.6 Experiment description
To assess the pattern detection feature of SmartGSN, we used SmartGSN to perform an experiment in which we ran each LLM five times using the aforementioned prompt. We performed that experiment five times to account for the non-deterministic nature of the LLMs SmartGSN uses.
4.1.7 Metrics used to assess experiment results
To assess the pattern detection results, we relied on well-known metrics [1]: Precision, Recall, and F-Measure. Thus, for a given assurance case, we compute the Precision as the number of patterns correctly detected by SmartGSN over the total number of patterns detected by SmartGSN. For a given assurance case, we compute the Recall as the number of patterns correctly detected by SmartGSN over the total number of patterns manually used to create that assurance case. Finally, we compute the F-Measure as the harmonic mean of the precision and the recall. Thus, we compute F-Measure as : (2 × Precision × Recall) / (Precision+Recall).
4.2 Results analysis: using SmartGSN for pattern detection
4.2.1 (RQ1) SmartGSN ability to detect patterns
Table 2 reports the average of recall (R), precision (P), and F-Measure (FM) for the five systems in our dataset using two models across various thresholds. We compute each average over five runs. The "Model" header either denotes GPT-4o or GPT-4 Turbo.
Table 2: Recall (R), Precision (P), and F-Measure (FM) results
System Model Metric Threshold
0.2 0.4 0.6 0.8 1
R P FM R P FM R P FM R P FM R P FM
ACAS XU GPT-4o 1 1 1 1 1 1 1 1 1 0 0 0 0 0 0
GPT-4 Turbo 1 1 1 1 1 1 1 1 1 0 0 0 0 0 0
BLUEROV2 GPT-4o 0.5 1 0.67 0.5 1 0.67 0.5 1 0.67 0 0 0 0 0 0
GPT-4 Turbo 0.5 1 0.67 0.5 1 0.67 0.5 1 0.67 0 0 0 0 0 0
GPCA GPT-4o 1 1 1 1 1 1 0 0 0 0 0 0 0 0 0
GPT-4 Turbo 1 1 1 1 1 1 0 0 0 0 0 0 0 0 0
IM Software GPT-4o 1 1 1 0 0 0 0 0 0 0 0 0 0 0 0
GPT-4 Turbo 1 1 1 0 0 0 0 0 0 0 0 0 0 0 0
DeepMind GPT-4o 1 1 1 1 1 1 0 0 0 0 0 0 0 0 0
GPT-4 Turbo 1 1 1 1 1 1 0 0 0 0 0 0 0 0 0
@@ 4.2 Results analysis: using SmartGSN for pattern detection
As Table 2 shows, for each system under analysis, both SmartGSN models achieve perfect metric scores (i.e. are able to detect patterns) at a threshold of 0.2, except for BLUEROV2, which has an average recall of 0.5 and an average FM-Measure of 0.67. At a threshold of 0.4, ACAS XU, GPCA, and DeepMind maintain perfect metric scores, whereas the BLUEROV2 still yields the same results. The IM Software metric scores drop to zero at this threshold. At a threshold of 0.8 or 1, all systems and models achieve zero scores across all metrics, indicating an inability for SmartGSN to detect patterns within an assurance case at these higher thresholds. Note that the values of the metrics are lower for BLUEROV2 because the assurance case of that system is derived from two patterns and not one pattern as the assurance cases of the other four systems under analysis. Still, our LLM prompts currently allow detecting a single pattern at the time. We plan to address this issue in future work.
4.2.2 (RQ2) Impact of varying metrics thresholds on SmartGSN pattern detection capabilities
As Table 2 shows, the threshold values of both metrics affect the SmartGSN pattern detection ability; lower thresholds like 0.2 result in detecting more patterns, while at higher thresholds (e.g., 0.8), performance drastically declines. The optimal threshold range for effective detection of patterns within assurance cases in our dataset likely falls within the [0.2, 0.6[ interval. Conducting more experiments in future work should confirm that observation.
5 Conclusion and future work
@@
In this paper, we presented a novel tool called SmartGSN. The latter leverages large language models (LLMs) to support the semi-automatic management of assurance cases. Our empirical evaluation focused on the pattern detection feature of that tool and showed that it can effectively rely on metrics and LLMs to detect assurance case patterns in assurance cases.
@@
Future work will focus on refining our pattern detection approach by exploring the use of more advanced rules and/or pattern-matching algorithms. Future work will also focus on extending our tool to support the refactoring of assurance cases using assurance case patterns. In addition, future work will focus on implementing a collaboration feature that will support the concurrent edition of the same assurance case when using SmartGSN. We are also planning to carry out a user study to gauge the usability of our tool in various application domains.
{credits}
5.0.1 Acknowledgements
The authors would like to thank the Mitacs GRI program for funding this work and Kimya K. Shahandashti for suggesting the use of React Flow.
[…]
The authors have no competing interests to declare.